OpenStudy (anonymous):

*** i will give you a medal *** Alicia can row 6 miles downstream in the same time it takes her to row 4 miles upstream. She rows downstream 3 miles/hour faster than she rows upstream. Find Alicia’s rowing rate each way. Round your answers to the nearest tenth, if necessary.

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

let x be the speed of boat, y be the speed of stream Alicia can row 6 miles downstream in the same time it takes her to row 4 miles upstream. => 6/(x+y) = 4(x-y).............equation #1 She rows downstream 3 miles/hour faster than she rows upstream. =>(x+y)-(x-y) = 3 => y = 1.5................#2 solving, x = 7.5 so, x+y = 9 x-y = 6 D- 9 mi/h downstream, 6 mi/h upstream